# Auto-Adaptive PINNs with Applications to Phase Transitions

### v1

@@ 1. Introduction
The study of Physics-Informed Neural Networks (PINNs) has grown rapidly in the past several years [1, 2]. One key issue in the study of these objects is the difference between applications to stationary and time-dependent problems. Despite theoretical similarities, consistent methods for the training of PINNs on time dependent problems remain elusive. We propose that choosing efficient and intelligent sampling distributions are potentially a key to alleviating these issues.
In the simulation of statics with PINNs it is a common technique to oversample, undersample, or otherwise highlight in regions which are known to be problematic such as regions containing shocks [3, 4], difficult boundary conditions [5], or other irregularities [6]. This is similar in spirit to multigrid schemes used extensively in traditional finite element methods, reviewed in [7]. In this context it is well understood that choosing a proper mesh can make the difference between a problem being completely infeasible and easily solvable. Generally in static PINNs, problem regions are identified post-hoc and then manually sampled more heavily. This allows for the gradual improvement of the static function over the network training. As the training continues it is paused periodically, at which point the sampling distributions can be edited as needed. We refer to this methodology as post-hoc sampling, as it done manually after some amount of training or detailed analysis of the particular approximation is done.
This post-hoc sampling is built on the fact that it is common in complex problems that some regions in the domain are more difficult to simulate than others. Simpler regions may have well-bounded error even without exceptionally low residual value, while other regions can explode in error with even the slightest inaccuracy. This is known as the conditioning of a problem [8, 9]. If small residual error leads to small actual accuracy, the problem is called well-conditioned. If a small residual does not necessarily result in small accuracy, the problem is called ill or poorly conditioned. More precisely, one can often bound the actual accuracy by some power of the loss times a constant. This constant is called the condition number. If it is very large, then the problem is poorly conditioned.
@@
To alleviate these issues, several methods have been experimented with. The first of these is Extended PINNs or XPINNs, first presented in [10]. This method involves the decomposition of the domain into non-overlapping subdomains. A separate PINN is then trained on each of these subdomains with an additional interior boundary condition guaranteeing that the PINNs match within the domain. Additional work has been done with this method to extend to variational problems [11] and to determine which problems the method is effective in solving [12]. In a similar spirit [13] proposes the partitioning of the time variable in their time slicing II approach, which will be a focus of our study.
@@
Another method designed to address these issues is the residual adaptive (or loss adaptive) PINN. Originally introduced in [14], this method is explored thoroughly in [15]. This involves the training of a PINN by constantly resampling according to the pointwise residual of the network. The precise nature of how this sampling is done varies, but the key idea is that the loss is reduced more effectively by highlighting regions of greater loss by sampling them more heavily. This method is extremely widely applied, but has the theoretical shortcoming that it assumes a uniform loss is optimal. For problems which are highly spatially irregular, or have regions where error grows at vastly different rates, these problems become more obvious. Of particular interest here is that the method is easily applied without prior knowledge of where the problematic regions will be, a strength that we hope to maintain in our own methodology.
@@
Our test case for these methods is the Allen-Cahn equation, which describes the phase separation of two fluids over time [16]. It does this through the evolution of the order parameter, which indicates the relative concentration of the two mixed fluids. We discuss these equations in more detail in section 1.2.2. In particular we experiment with an energy adaptive variant of the described auto adaptive sampling, where we sample in proportion to the pointwise energy density of the Allen-Cahn equation. As we show heuristically in 1.2.3, regions of high energy often correspond well with the problematic regions described earlier. These regions contain behavior where a small residual can invite a disproportionate amount of error. Additionally these regions are often areas more important to the continued accuracy as time is evolved.
This setting highlights the issues that we hope to address with our adaptive methods. Allen-Cahn displays a separation of time and spatial scales as the interfaces quickly form and slowly move. Though these interfaces only partially describe where problematic regions occur. In particular, we contrast our results with those of [13], who also experiment with Allen-Cahn and the related Cahn-Hilliard model. The authors pioneer the time-slicing method by testing on the Allen-Cahn and Cahn-Hilliard systems. In their experiments, they encounter several issues that we hope to address. First, they observe consistent large error on the interfaces formed in the Allen-Cahn evolution as well as other regions where the function is away from $\pm 1$. We will show that the energy adaptive approach greatly alleviates this issue. Additionally for problems which have particularly large separation of scales, their “time-adaptive I” approach fails and they are forced to instead train several networks on the same problem in what they call the “Time-Adaptive II” approach. We find this solution undesirable, as the training of multiple networks may be expensive.
@@ 1.1. Physics Informed Neural Networks
First introduced in [17], a Physics-Informed Neural Network is a method of approximating a partial differential equation (PDE) using a Neural Network structure with artificial ‘data’ derived from the PDE. In this framework, we define a neural network which is itself an approximation to the solution to a PDE. We then iteratively update this approximation through gradient descent on a loss function, which measures the fidelity of the approximation to the governing equations.
@@ 1.2.2. Allen-Cahn Equation
This has been shown as an appropriate choice of $\Psi$, though sometimes truncations are used to avoid numerical difficulties [18]. This yields the following final equation, which amount to a full description of our problem.
@@ 2.1. Metropolis-Hastings Algorithm
In order to efficiently sample from a probability distribution dependent on the network and its derivatives, we use the Metropolis-Hastings Algorithm [19, 20]. This is a sampling method which iteratively improves a collection of randomly chosen points to match the distribution given by a known probability density function. This method avoids the need to invert a distribution, and is thus ideal for settings where direct sampling is difficult.
@@
It is worth noting that there is a brief history of the use of Metropolis-Hastings and the broader class of Markov-Chain-Monte-Carlo (MCMC) methods in PINNs. In particular, the results of [15] suggest that this is the most effective method for the implementation of residual adaptive sampling. Though dropout methods remain more popular in residual adaptive methods, such as was employed in [13], because of their ease of implementation. Additionally, the Bayesian-PINN (B-PINN) [21] uses similar MCMC-based inference to capture uncertainty in the network parameters, addressing noise through probabilistic modeling rather than deterministic regularization.
The Metropolis-Hastings is additionally very well-studied and lends itself very naturally to the context of Physics-Informed Neural Networks. As an iteration based method, it can be run easily in conjunction with training. Additional improvements such as multiple proposal [22] can nicely address common issues in PDEs and in shocks in particular. Additionally, the method is nicely parrallelizable and can be run very efficiently on GPUs [23]. These benefits make it an exceptionally good fit for sampling in the context of Neural Networks.
@@ 3.1. Time Slicing
This method was introduced in [13] and involves the gradual expansion of the time domain in discrete steps. The idea is to reduce the time complexity of the problem by first greatly restricting the time domain until it is palatable to the network. Once the problem is learned on that smaller interval, the interval can be increased. This allows for the network to only require learning a small time interval at any given time time, which is generally easier for training. This method was tested very thoroughly with a large degree of success. However, for higher-dimensional and less regular systems further issues arise.
@@
The authors of [13] were aware of this issue, so they presented an alternative method that they call ‘Time Slicing II.’ This method involves the training of a separate network on each time slice rather than using the gradual expansion of training time on a single network. This does resolve the issue of catastrophic unlearning, since no network is trained to perform multiple tasks. However, it is somewhat undesirable as it requires many networks, a true time discretization, and can also dramatically increase cost in training time and number of network parameters. For these reasons we manually edit the learning rate as we evolve through time slices, which is discussed next.
@@ 3.2. Learning Rate Schedule
We additionally use a learning rate scheduler, which updates the learning rate as the training process progresses. We use a fairly simple scheduler which updates the learning rate only based off the current time slice being trained. This helps to combat the catastrophic unlearning described in the section above. We decrease the learning rate roughly linearly as the progress through the time-slices. More sophisticated methods of choosing the learning rates are constantly being studied [24, 25], however for simplicity and ease of implementation we choose to use only to use this simple scheme.
@@ 3.3. Residual Adaptive Sampling
A commonly employed method to increase training efficacy is Residual Adaptive sampling, first proposed in [14]. This technique involves sampling the domain in proportion to the current pointwise value of the residual. This leads to decreasing the loss very efficiently, as has been thoroughly explored in [15], and seen in many other papers including [13]. In our numerical experiments, we will use this method as a test to compare our own methods against.
@@ 3.6. Latin Hypercube Sampling
For uniform sampling in our domain, we use Latin Hypercube sampling. This method reduces variance in sampling by partitioning the domain into a grid of $N^{d}$ hypercubes, where $N$ is the desired number of sample points and $d$ is the number of spatial dimensions. Then hypercubes are sampled randomly so that the selected hypercubes are orthogonal. This yields a selection of $N$ hypercubes in the grid, with no selected hypercube in the same row or column as any other. Once this orthogonal set of hypercubes is selected, one point within each is chosen randomly from a uniform distribution on that hypercube. These individual sample points combine to form a representative sample of $N$ points from the domain. Further explanation can be found [26]. This method guarantees that the uniform sample will be representative of the entire domain. This sampling also greatly reduces variance, and thus is less reliant on the convergence yielded by the law of large numbers. This especially helps to reduce inconsistencies of sampling too few points, and makes the training more robust.
@@ 3.8. Optimizer Choice
Answering these questions amounts to numerical experiment. We found that the implementation of BFGS at the conclusion of each time slice hurt overall results. We speculate that the method did too good a job at ‘locking in’ the behavior of the system early on, which prevented the network from adapting from the fast behavior in the first portion of the time interval to the slow behavior in the later portion of the interval. For this reason we apply BFGS only at the conclusion of the full time interval’s training. Notably, this is not how the method was implemented in [13]. However, they provide no discussion of this decision.
@@ 3.9. Algorithm Schematic
Algorithm 1 PINN Training with Adaptive Collocation
@@ 4. Numerical Results
The following numerical benchmarks are taken from [13]. In the examples they provide, there is a clear concentration of the error in the regions of high energy. As such, this is an ideal setting for the testing of the improvements given by our new method.
For each example, we compare three methods. First, we run a high fidelity finite difference to yield an ‘exact’ solution. The second method is Residual Adaptive sampling. Here we use the Metropolis Hastings Algorithm to sample in proportion to the loss and combine these with points sampled uniformly as in [15]. Additionally this implementation uses all of the techniques described in the implementation section, including time slices, learning rate scheduling, minibatching, latin hypercube sampling, and initial condition weights. This is meant to approximate current state of the art techniques. Finally we implement our energy adaptive method described in section 3.
All simulations for this project are done in python using the PyTorch package for Neural Networks, and run on an nVIDIA Quatro RTX A6000 GPU, provided generously for use by the Institute for Scientific Computing and Applied Mathematics at Indiana University. All scripts can be found online at https://github.com/kevmbuck/Energy-Adaptive-PINNS.
@@
We use experimental hyperparameters all identical to the first example, in order to test the robustness of each method without fine tuning. The problems are also very similar, so these hyperparameters are likely not too far from optimal.
@@
Finally we experiment with the 2D Example given in [13]. This is characterized by the parameters
@@
We can see this difficulty manifest in Figure 7. Here we see that the expansion of the training time from $t=9$ to $t=10$ results in the catastrophic unlearning described in the introduction. Not only do we loose accuracy on the final time slice, but we also loose accuracy on the entire preceding time domain. This can also be seen in the decay of the loss in Figure 8. Here we see that the loss levels off despite the reduction in learning rate later in the process, indicating a problem in training.
@@
This problem accentuates the issues present in time-dependent simulation of multiscale systems. While we believe that time-dependent adaptive sampling could result in a resolution of this issue, we leave it as future work for now.
@@ 5. Conclusions and Future Directions
Analytic verification in the form of a proof is also desirable for the energy adaptive method, perhaps in a similar manner to the proofs in [27]. If well understood, this could also motivate choices of sampling densities for other difficult problems.

### v3

@@ 1. Introduction
The study of Physics-Informed Neural Networks (PINNs) has grown rapidly in the past several years [review, review2]. One key issue in the study of these objects is the difference between applications to stationary and time-dependent problems. Despite theoretical similarities, consistent methods for the training of PINNs on time dependent problems remain elusive. We propose that choosing efficient and intelligent sampling distributions are potentially a key to alleviating these issues.
In the simulation of statics with PINNs it is a common technique to oversample, undersample, or otherwise highlight in regions which are known to be problematic such as regions containing shocks [Shock1, Shock2], difficult boundary conditions [BoundaryPINN], or other irregularities [Javi]. This is similar in spirit to multigrid schemes used extensively in traditional finite element methods, reviewed in [multiscalerev]. In this context it is well understood that choosing a proper mesh can make the difference between a problem being completely infeasible and easily solvable. Generally in static PINNs, problem regions are identified post-hoc and then manually sampled more heavily. This allows for the gradual improvement of the static function over the network training. As the training continues it is paused periodically, at which point the sampling distributions can be edited as needed. We refer to this methodology as post-hoc sampling, as it done manually after some amount of training or detailed analysis of the particular approximation is done.
This post-hoc sampling is built on the fact that it is common in complex problems that some regions in the domain are more difficult to simulate than others. Simpler regions may have well-bounded error even without exceptionally low residual value, while other regions can explode in error with even the slightest inaccuracy. This is known as the conditioning of a problem [NLA, me]. If small residual error leads to small actual accuracy, the problem is called well-conditioned. If a small residual does not necessarily result in small accuracy, the problem is called ill or poorly conditioned. More precisely, one can often bound the actual accuracy by some power of the loss times a constant. This constant is called the condition number. If it is very large, then the problem is poorly conditioned.
@@
To alleviate these issues, several methods have been experimented with. The first of these is Extended PINNs or XPINNs, first presented in [xPINN]. This method involves the decomposition of the domain into non-overlapping subdomains. A separate PINN is then trained on each of these subdomains with an additional interior boundary condition guaranteeing that the PINNs match within the domain. Additional work has been done with this method to extend to variational problems [varxPINN] and to determine which problems the method is effective in solving [whenxPINN]. In a similar spirit [WZ] proposes the partitioning of the time variable in their time slicing II approach, which will be a focus of our study.
@@
Another method designed to address these issues is the residual adaptive (or loss adaptive) PINN. Originally introduced in [orig_resadapt], this method is explored thoroughly in [resadapt]. This involves the training of a PINN by constantly resampling according to the pointwise residual of the network. The precise nature of how this sampling is done varies, but the key idea is that the loss is reduced more effectively by highlighting regions of greater loss by sampling them more heavily. This method is extremely widely applied, but has the theoretical shortcoming that it assumes a uniform loss is optimal. For problems which are highly spatially irregular, or have regions where error grows at vastly different rates, these problems become more obvious. Of particular interest here is that the method is easily applied without prior knowledge of where the problematic regions will be, a strength that we hope to maintain in our own methodology.
@@
Our test case for these methods is the Allen-Cahn equation, which describes the phase separation of two fluids over time [AC]. It does this through the evolution of the order parameter, which indicates the relative concentration of the two mixed fluids. We discuss these equations in more detail in section 1.2.2. In particular we experiment with an energy adaptive variant of the described auto adaptive sampling, where we sample in proportion to the pointwise energy density of the Allen-Cahn equation. As we show heuristically in 1.2.3, regions of high energy often correspond well with the problematic regions described earlier. These regions contain behavior where a small residual can invite a disproportionate amount of error. Additionally these regions are often areas more important to the continued accuracy as time is evolved.
This setting highlights the issues that we hope to address with our adaptive methods. Allen-Cahn displays a separation of time and spatial scales as the interfaces quickly form and slowly move. Though these interfaces only partially describe where problematic regions occur. In particular, we contrast our results with those of [WZ], who also experiment with Allen-Cahn and the related Cahn-Hilliard model. The authors pioneer the time-slicing method by testing on the Allen-Cahn and Cahn-Hilliard systems. In their experiments, they encounter several issues that we hope to address. First, they observe consistent large error on the interfaces formed in the Allen-Cahn evolution as well as other regions where the function is away from $\pm 1$. We will show that the energy adaptive approach greatly alleviates this issue. Additionally for problems which have particularly large separation of scales, their “time-adaptive I” approach fails and they are forced to instead train several networks on the same problem in what they call the “Time-Adaptive II” approach. We find this solution undesirable, as the training of multiple networks may be expensive.
@@ 1.1. Physics Informed Neural Networks
First introduced in [PINN], a Physics-Informed Neural Network is a method of approximating a partial differential equation (PDE) using a Neural Network structure with artificial ‘data’ derived from the PDE. In this framework, we define a neural network which is itself an approximation to the solution to a PDE. We then iteratively update this approximation through gradient descent on a loss function, which measures the fidelity of the approximation to the governing equations.
@@ 1.2.2. Allen-Cahn Equation
This has been shown as an appropriate choice of $\Psi$, though sometimes truncations are used to avoid numerical difficulties [GinzLandApprox]. This yields the following final equation, which amount to a full description of our problem.
@@ 2.1. Metropolis-Hastings Algorithm
In order to efficiently sample from a probability distribution dependent on the network and its derivatives, we use the Metropolis-Hastings Algorithm [Metropolis, Hastings]. This is a sampling method which iteratively improves a collection of randomly chosen points to match the distribution given by a known probability density function. This method avoids the need to invert a distribution, and is thus ideal for settings where direct sampling is difficult.
@@
It is worth noting that there is a brief history of the use of Metropolis-Hastings and the broader class of Markov-Chain-Monte-Carlo (MCMC) methods in PINNs. In particular, the results of [resadapt] suggest that this is the most effective method for the implementation of residual adaptive sampling. Though dropout methods remain more popular in residual adaptive methods, such as was employed in [WZ], because of their ease of implementation. Additionally, the Bayesian-PINN (B-PINN) [BPINN] uses similar MCMC-based inference to capture uncertainty in the network parameters, addressing noise through probabilistic modeling rather than deterministic regularization.
The Metropolis-Hastings is additionally very well-studied and lends itself very naturally to the context of Physics-Informed Neural Networks. As an iteration based method, it can be run easily in conjunction with training. Additional improvements such as multiple proposal [MultiMCMC] can nicely address common issues in PDEs and in shocks in particular. Additionally, the method is nicely parrallelizable and can be run very efficiently on GPUs [ParMCMC]. These benefits make it an exceptionally good fit for sampling in the context of Neural Networks.
@@ 3.1. Time Slicing
This method was introduced in [WZ] and involves the gradual expansion of the time domain in discrete steps. The idea is to reduce the time complexity of the problem by first greatly restricting the time domain until it is palatable to the network. Once the problem is learned on that smaller interval, the interval can be increased. This allows for the network to only require learning a small time interval at any given time time, which is generally easier for training. This method was tested very thoroughly with a large degree of success. However, for higher-dimensional and less regular systems further issues arise.
@@
The authors of [WZ] were aware of this issue, so they presented an alternative method that they call ‘Time Slicing II.’ This method involves the training of a separate network on each time slice rather than using the gradual expansion of training time on a single network. This does resolve the issue of catastrophic unlearning, since no network is trained to perform multiple tasks. However, it is somewhat undesirable as it requires many networks, a true time discretization, and can also dramatically increase cost in training time and number of network parameters. For these reasons we manually edit the learning rate as we evolve through time slices, which is discussed next.
@@ 3.2. Learning Rate Schedule
We additionally use a learning rate scheduler, which updates the learning rate as the training process progresses. We use a fairly simple scheduler which updates the learning rate only based off the current time slice being trained. This helps to combat the catastrophic unlearning described in the section above. We decrease the learning rate roughly linearly as the progress through the time-slices. More sophisticated methods of choosing the learning rates are constantly being studied [lrannealing, softadapt], however for simplicity and ease of implementation we choose to use only to use this simple scheme.
@@ 3.3. Residual Adaptive Sampling
A commonly employed method to increase training efficacy is Residual Adaptive sampling, first proposed in [orig_resadapt]. This technique involves sampling the domain in proportion to the current pointwise value of the residual. This leads to decreasing the loss very efficiently, as has been thoroughly explored in [resadapt], and seen in many other papers including [WZ]. In our numerical experiments, we will use this method as a test to compare our own methods against.
@@ 3.6. Latin Hypercube Sampling
For uniform sampling in our domain, we use Latin Hypercube sampling. This method reduces variance in sampling by partitioning the domain into a grid of $N^{d}$ hypercubes, where $N$ is the desired number of sample points and $d$ is the number of spatial dimensions. Then hypercubes are sampled randomly so that the selected hypercubes are orthogonal. This yields a selection of $N$ hypercubes in the grid, with no selected hypercube in the same row or column as any other. Once this orthogonal set of hypercubes is selected, one point within each is chosen randomly from a uniform distribution on that hypercube. These individual sample points combine to form a representative sample of $N$ points from the domain. Further explanation can be found [hypercube]. This method guarantees that the uniform sample will be representative of the entire domain. This sampling also greatly reduces variance, and thus is less reliant on the convergence yielded by the law of large numbers. This especially helps to reduce inconsistencies of sampling too few points, and makes the training more robust.
@@ 3.8. Optimizer Choice
Answering these questions amounts to numerical experiment. We found that the implementation of BFGS at the conclusion of each time slice hurt overall results. We speculate that the method did too good a job at ‘locking in’ the behavior of the system early on, which prevented the network from adapting from the fast behavior in the first portion of the time interval to the slow behavior in the later portion of the interval. For this reason we only apply ADAM for our results.
@@ 3.9. Algorithm Schematic
Algorithm 1 PINN Training with Auto-Adaptive Sampling
@@ 4. Numerical Results
The following numerical benchmarks are taken from [WZ]. In the examples they provide, there is a clear concentration of the error in the regions of high energy. As such, this is an ideal setting for the testing of the improvements given by our new method.
For each example, we compare three methods. First, we run a high fidelity finite difference to yield an ‘exact’ solution. The second method is Residual Adaptive sampling. Here we use the Metropolis Hastings Algorithm to sample in proportion to the loss and combine these with points sampled uniformly as in [resadapt]. Additionally this implementation uses all of the techniques described in the implementation section, including time slices, learning rate scheduling, minibatching, latin hypercube sampling, and initial condition weights. This is meant to approximate current state of the art techniques. Finally we implement our energy adaptive method described in section 3.
All simulations for this project are done in python using the PyTorch package for Neural Networks, and run on an NVIDIA Quatro RTX A6000 GPU, provided generously for use by the Institute for Scientific Computing and Applied Mathematics at Indiana University. All scripts can be found online at https://github.com/kevmbuck/Auto-Adaptive-PINNs.
@@
We use experimental hyperparameters all identical to the first example, in order to test the robustness of each method without fine tuning. The problems are also very similar, so these hyperparameters are not too far from optimal.
@@
Finally we experiment with the 2D Example given in [WZ]. This is characterized by the parameters
@@
We can see this difficulty manifest in Figure 7, which depicts results from the baseline residual adaptive method. Here we see that the expansion of the training time from $t=9$ to $t=10$ results in the catastrophic unlearning described in the introduction. Not only do we loose accuracy on the final time slice, but we also loose accuracy on the entire preceding time domain.
We next show the results from energy adaptive training in Figure 8. We see here that the errors are much smaller and more contained than in the residual adaptive method. The precise error values can additionally be found in Table 3. We see that the energy adaptive method when trained until time $9$ leads to very good extrapolation until time $10$. Here the errors are acceptable even for a simulation on the full domain. Despite the improved accuracy, the energy adaptive method still succumbs to the catastrophic unlearning phenomenon in the final time slice. Although even after the unlearning the errors are still comparable to the residual adaptive method before unlearning.
@@
Error Measure / Method Trained until Residual Adaptive Energy Adaptive
Relative $L^{2}$ at $T=10$ $T=9$ 1.52e-01 3.56e-02
Relative $L^{2}$ at $T=10$ $T=10$ 3.53e-01 1.93e-01
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=9$ 1.73e+$0$ 9.79e-01
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=10$ 1.98e+$0$ 1.90e+$0$
Table 3. We observe the errors of each method.
In general there are several methods which can be used to reduce the problems posed by catastrophic unlearning. The two most prominent of which would be the employment of an xPINN [xPINN] or the use of exponential time sampling [ExpTime]. These directions are known to be effective, but require a post-hoc knowledge of where the unlearning occurs. For our analysis here, we are trying to confine our approach only to methods which can be applied without prior knowledge of the location of the unlearning. We also believe that an intelligently chosen heuristic employed via auto-adaptive PINNs could efficiently resolve this issue. However, we leave the exploration of this as a future direction.
@@ 5. Conclusions and Future Directions
Analytic verification in the form of a proof is also desirable for the energy adaptive method, perhaps in a similar manner to the proofs in [ExpTime]. If well understood, this could also motivate choices of sampling densities for other difficult problems.
@@
Acknowledgments
We would like to thank the Institute of Scientific Computing and Applied Mathematics for generously providing access to their computing resources, including the two NNVIDIA A6000 GPUs used for our computations. The work of author Kevin Buck was also partially funded by NSF Grants number DMS-2154387 and DMS-2206105.